OpenStudy (anonymous):

one angle of an isosceles triangle is 34 degrees. What are the other 2 angles 73 and 34 73 and 73 34 and 112 34 and 152 there is more then one answerer

ganeshie8 (ganeshie8):

suppose the given angle is the vertex angle, then the base angles would be congruent : `34 + x + x = 180` solve x

ganeshie8 (ganeshie8):

that gives you one solution ^^

OpenStudy (anonymous):

number2, 3

ganeshie8 (ganeshie8):

what do u mean

ganeshie8 (ganeshie8):

34 + x + x = 180 34 + 2x = 180 2x = 146 x = 73

ganeshie8 (ganeshie8):

so answer is either A or B

OpenStudy (anonymous):

so one would be 73+73+64 would be one right

ganeshie8 (ganeshie8):

Oh yes, you're right !!!!! there will be one more solution

ganeshie8 (ganeshie8):

if 34 is the base angle : `34 + 34 + x = 180` solve x

OpenStudy (anonymous):

34+112+34

ganeshie8 (ganeshie8):

so below are the answers is it ``` 73 and 73 34 and 112 ```

ganeshie8 (ganeshie8):

nice :)